Tuba for the month of August, had sales of $800,000; variable costs of 60% of sales; and total fixed costs of $215,000. Determine the dollar sales needed to achieve a target income of $200,000 in September. (Ignore taxes)
Answer #1

Contribution margin rate = 1 - 0.6 = 40%

Dollar sales needed to achieve target profit = (Fixed asset + Target profit) / Contribution margin rate

= (215000+200000) /40% = $1, 037,500